DeAnne Company produces a single product. The company's variable costing income statement for August appears below: DeAnne Company Income Statment For the month ended August 31 Sales ($20 per unit) $ 848,000 Variable expenses: Variable cost of goods sold 508,800 Variable selling expense 84,800 Total variable expenses 593,600 Contribution margin 254,400 Fixed expenses: Fixed manufacturing overhead 141,600 Fixed selling and administrative 70,800 Total fixed expenses 212,400 Net operating income $ 42,000 The company produced 35,400 units in August and the beginning inventory consisted of 8,020 units. Variable production costs per unit and total fixed costs have remained constant over the past several months. The value of the company's inventory on August 31 under absorption costing would be: $16,320 $18,360 $23,360 $12,240.
Answer #1

First lets calculate the closing inventory in units

Closing inventory = Opening inventory + Units produced - Units sold

Units sold = sales value / selling price per unit

= $848000/$20 = 42400

Closing inventory in units = 8020+35400-42400 = 1020

Now to arrive at the value of the units costs upto production will be allocted to the units produced

here selling expenses are excluded

So the expenses are

Variable cost of goods sold = $508800

Units sold = 42400

Variable cost per unit = $508800/42400 = $12

Fixed manufacturing overhead = $141600

Units Produced = 35400

So Fixed manufacturing overhead per unit = $141600+/35400 = $4

So the Value of one Unit = $12+$4 = $16

So the value of Closing stock = $16*1020 = $16320

So the correct answer is first option $16320

Please free to comment for further clarification and upvote if it was helpful